# Principled Federated Random Forests for Heterogeneous Data

### v1

@@ 6 Conclusion
We propose FedForest, a federated forests procedure with formal guarantees, proving that centralized CART and RF behavior can be exactly recovered from federated data. Our approach serves as a foundational building block for extending the entire family of CART-criterion-based methods, including causal, quantile, and survival Forests, to heterogeneous federated environments.
Impact Statement
This paper presents work whose goal is to advance the field of machine learning. There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
Appendix A Proofs
A.1 Impurity Decomposition
Proof of Theorem 3.3 which states:
$I(\mathcal{S}_{\nu})=\sum_{k=1}^{K}\frac{n_{\nu,k}}{n_{\nu}}I(\mathcal{S}_{\nu,k})+\mathcal{E}(\nu),$
where the heterogeneity term $\mathcal{E}(\nu)$ is:
where $\hat{P}_{k}$ is the empirical distribution of client $k$’s outcome and $\Psi$ is the impurity function from statistics, and for variance, Gini and entropy impurity functions we have $\mathcal{E}(\nu)\geq 0$ and:
$\mathcal{E}(\nu)=0\iff\forall(k,k^{\prime}),\mathbb{E}[Y\mid X,\mathcal{S}_{\nu},H=k]=\mathbb{E}[Y\mid X,\mathcal{S}_{\nu},H=k^{\prime}].$
Proof.
Let $\mathcal{S}_{\nu}$ be the set of samples at node $\nu$, and $\mathcal{S}_{\nu,k}$ the subset of samples belonging to client $k$, such that $\mathcal{S}_{\nu}=\bigcup_{k=1}^{K}\mathcal{S}_{\nu,k}$ and $\mathcal{S}_{\nu,k}\cap\mathcal{S}_{A,k^{\prime}}=\emptyset$. Let $P_{\nu}$ denote the empirical distribution of the target variable $Y$ over the pooled sample $\mathcal{S}_{\nu}$, and $P_{\nu,k}$ denote the empirical distribution over the local sample $\mathcal{S}_{\nu,k}$. The empirical distribution is defined as the measure assigning mass $1/n$ to each observation. Since the datasets are disjoint, the global empirical measure $P_{\nu}$ is the weighted average of the local measures. For any event $E$:
$P_{\nu}(E)$ $=\frac{1}{n_{\nu}}\sum_{i\in\mathcal{S}_{\nu}}\mathbb{1}(y_{i}\in E)$
$=\sum_{k=1}^{K}w_{\nu,k}P_{\nu,k}(E),$
The global empirical impurity is defined as the function $\Psi$ (mean squared error, Gini, cross entropy etc.) applied to the global distribution $P_{\nu}$:
$I(\mathcal{S}_{\nu})$ $=\Psi\left(\sum_{k=1}^{K}w_{\nu,k}P_{\nu,k}\right)$
$=\underbrace{\sum_{k=1}^{K}w_{\nu,k}\Psi(P_{\nu,k})}_{\text{Weighted Local Impurities}}+\underbrace{\left[\Psi\left(\sum_{k=1}^{K}w_{\nu,k}P_{\nu,k}\right)-\sum_{k=1}^{K}w_{\nu,k}\Psi(P_{\nu,k})\right]}_{\text{Heterogeneity Term }}$
$=\sum_{k=1}^{K}\frac{n_{\nu,k}}{n_{\nu}}I(\mathcal{S}_{\nu,k})+\mathcal{E}(\nu),$
which corresponds to (6).
Finally, we show that for strictly concave impurities, $\mathcal{E}(\nu)=0\iff\forall(k,k^{\prime}),\hat{\mathbb{E}}[Y\mid\mathcal{S}_{\nu},H=k]=\hat{\mathbb{E}}[Y\mid\mathcal{S}_{\nu},H=k^{\prime}]$. Let $\mu_{\nu,k}=\hat{\mathbb{E}}[Y\mid\mathcal{S}_{\nu},H=k]$ denote the local empirical mean outcome (regression) or class probability vector (classification).
For classification, $\mu_{\nu,k}$ is the vector of class probabilities (since $\hat{\mathbb{E}}[Y_{\text{one-hot}}]=\mathbf{p}$), and impurity is defined as $\Psi(\mu_{\nu,k})$ where $\Psi$ is strictly concave (e.g., Gini or Entropy). By Jensen’s inequality, the gap $\mathcal{E}(\nu)=\Psi(\sum\frac{n_{\nu,k}}{n_{\nu}}\mu_{\nu,k})-\sum\frac{n_{\nu,k}}{n_{\nu}}\Psi(\mu_{\nu,k})$ is non-negative.
In both cases, due to strict concavity, $\mathcal{E}(\nu)=0\iff\mu_{\nu,1}=\dots=\mu_{\nu,K}$, i.e., the mean outcome is identical across sites. ∎
A.2 Impurity reduction estimator: homogeneous setting
Proof of Theorem 3.4:
Proof.
Recall the definition of the global centralized empirical impurity reduction at node $A$:
$\Delta I(j,t;\mathcal{S}_{\nu})=I(\mathcal{S}_{\nu})-\left[\frac{n_{L}}{n_{\nu}}I(\mathcal{S}_{L})+\frac{n_{R}}{n_{\nu}}I(\mathcal{S}_{R})\right].$
Using the impurity decomposition in Eq. (6), we have:
$I(\mathcal{S}_{\nu})$ $=\sum_{k=1}^{K}\frac{n_{\nu,k}}{n_{\nu}}I(\mathcal{S}_{\nu,k})+\mathcal{E}(A),$
$I(\mathcal{S}_{L})$ $=\sum_{k=1}^{K}\frac{n_{L,k}}{n_{L}}I(\mathcal{S}_{L,k})+\mathcal{E}(L),$
$I(\mathcal{S}_{R})$ $=\sum_{k=1}^{K}\frac{n_{R,k}}{n_{R}}I(\mathcal{S}_{R,k})+\mathcal{E}(R),$
which yields:
$\Delta I(j,t;\mathcal{S}_{\nu})=\underbrace{\sum_{k=1}^{K}\frac{n_{\nu,k}}{n_{\nu}}\Delta I_{k}(j,t;\mathcal{S}_{\nu,k})}_{\widehat{\Delta I}_{\text{fed}}}+\underbrace{\left(\mathcal{E}(A)-\frac{n_{L}}{n_{\nu}}\mathcal{E}(L)-\frac{n_{R}}{n_{\nu}}\mathcal{E}(R)\right)}_{\text{Residual Term }\delta_{n}}.$
Thus, the discrepancy is strictly governed by the heterogeneity terms: $|\widehat{\Delta I}_{\text{fed}}-\Delta I|=|\delta_{n}|$.
We analyze $\delta_{n}$ for the standard impurity measures derived in the previous section.
Variance Impurity (Regression). When impurity is measured as variance, the heterogeneity term $\mathcal{E}$ is explicitly the variance of the local means:
$\mathcal{E}(\nu)=\sum_{k=1}^{K}\frac{n_{\nu,k}}{n_{\nu}}(\bar{y}_{\nu,k}-\bar{y}_{\nu})^{2}.$
Under 2.1, each client draws samples from the same population. At node $A$, the local sample mean at node $\bar{y}_{\nu,k}$ is an unbiased estimator of $\mu_{\nu}$ with variance $\mathbb{V}(\bar{y}_{\nu,k})=\frac{\sigma_{\nu}^{2}}{n_{\nu,k}}$. The term $(\bar{y}_{\nu,k}-\bar{y}_{\nu})^{2}$ scales with its variance at $O_{p}(1/n_{\nu,k})$:
$\mathcal{E}(\nu)=\sum_{k=1}^{K}\frac{n_{\nu,k}}{n_{\nu}}O_{p}\left(\frac{1}{n_{\nu,k}}\right)=\sum_{k=1}^{K}O_{p}\left(\frac{1}{n_{\nu}}\right)=O_{p}\left(\frac{K}{n_{\nu}}\right).$
Gini and entropy impurities (classification). In classification, the local ”mean” corresponds to the vector of class probabilities $\mathbf{p}_{\nu,k}$. For Gini impurity, the heterogeneity term is exactly the weighted squared Euclidean distance between the local and global probability vectors (by the law of total variance):
$\mathcal{E}_{\text{Gini}}(\nu)=\sum_{k=1}^{K}\frac{n_{\nu,k}}{n_{\nu}}\|\mathbf{p}_{\nu,k}-\mathbf{p}_{\nu}\|^{2}.$
For Entropy, the heterogeneity corresponds to the Jensen-Shannon divergence, which acts locally as a quadratic approximation. By a second-order Taylor expansion, it relates to the chi-squared distance:
$\mathcal{E}_{\text{Entropy}}(A)=\sum_{k=1}^{K}\frac{n_{\nu,k}}{n_{\nu}}D_{\text{KL}}(\mathbf{p}_{\nu,k}||\mathbf{p}_{\nu})\approx\frac{1}{2}\sum_{k=1}^{K}\frac{n_{\nu,k}}{n_{\nu}}\sum_{c=1}^{C}\frac{(p_{\nu,k}^{(c)}-p_{\nu}^{(c)})^{2}}{p_{\nu}^{(c)}}.$
In both cases, the heterogeneity is driven by the squared norm $\|\mathbf{p}_{\nu,k}-\mathbf{p}_{\nu}\|^{2}$. Since the empirical class frequencies $\mathbf{p}_{\nu,k}$ deviate from the true population probabilities by a factor of $O_{p}(1/\sqrt{n_{\nu,k}})$, the squared norm scales as $O_{p}(1/n_{\nu,k})$.
To conclude, since $\mathcal{E}(L)$ and $\mathcal{E}(R)$ follow the same scaling (with denominators $n_{L},n_{R}<n_{\nu}$), the total residual $\delta_{n}$ is a linear combination of terms that are all $O_{p}(K/n_{\nu})$. Therefore:
$\left|\Delta I_{n}(j,t;\mathcal{S}_{\nu})-\sum_{k=1}^{K}\frac{n_{\nu,k}}{n_{\nu}}\Delta I_{n_{k}}(j,t;\mathcal{S}_{\nu,k})\right|=O_{p}\left(\frac{K}{n_{\nu}}\right).$
A.3 Candidate thresholds with quantile sketching
Proof of Theorem 3.1:
Proof.

### v2

@@ 6 Conclusion
We propose FedForest, a federated forest training procedure with formal guarantees, showing that centralized CART and Random Forest behavior can be faithfully recovered from decentralized data without sharing individual samples. By reconstructing global split criteria from local summaries, our approach provides a principled foundation for extending the broader family of CART-based methods—including causal, quantile, and survival forests—to heterogeneous federated settings.
While exchanging only summary statistics reduces direct privacy exposure, these aggregates may still leak sensitive information about local data. Our protocol can be further strengthened with additional safeguards, including structural regularization (e.g., depth constraints and minimum leaf sizes), adaptive binning ($B_{\nu}<\min_{k}n_{\nu,k}$), secure aggregation, and client abstention for small local partitions. More broadly, an interesting direction for future work is to extend our framework with formal differential privacy guarantees and to study the resulting privacy–utility trade-offs.
Appendix A FedForest Communication and Computation Costs
In this section, we detail the communication and computation complexities of our proposed FedForest protocol and compare them against standard baselines. Let $T$ be the number of trees, $M$ the maximum tree depth, $d$ the number of features, $K$ the number of clients, $n_{k}$ the number of samples at client $k$, and $n$ the total number of samples.
Communication per node.
We count communication in transmitted scalars for one node $\nu$ and one client $k$, with $B$ representing the quantile levels (sketch size), and $S$ representing the sufficient-statistic size (e.g., $S=3$ for regression and $S=C$ for $C$-class classification). All costs below are strictly independent of the node sample sizes $(n_{\nu,k},n_{\nu})$.
  • FedForest (exact gain): Communication occurs in two steps. (i) Sketching: the client sends $d$ univariate quantile sketches of size $B$, for a cost of $dB$. (ii) Evaluation: for each tested threshold in $\mathcal{T}$, the client returns the left-child sufficient statistics, costing $S$ scalars per candidate. The total cost of FedForest per node is therefore $O(dBS)$.
  • Top-$L$ shortlisting: If the server evaluates only $|\mathcal{L}_{\nu}|=L\ll d$ shortlisted features, the evaluation term strictly reduces to $O(LBS)$.
  • AvgImp (Top-$L$ + averaged local gains): If clients compute their impurity gains locally at the candidate thresholds of the shortlisted features, they send one scalar per threshold instead of $S$. Combining Top-$L$ with AvgImp therefore reduces the evaluation communication to $O(LB)$, at the price of shifting computation from the server to the clients for the local gain evaluation.
Computation complexity.
In a centralized Random Forest, all computation is localized at the server, requiring $O(T2^{M}dn)$ operations to route data and evaluate splits across all nodes. In FedForest, this burden is decoupled:
  • Client-side computation: At each level of the tree, a client must route its local data and update the quantile sketches for the active nodes. Over $T$ trees and $M$ depth levels, this requires $O(TMdn_{k})$ operations, scaling only with the client’s local dataset size. For the AvgImp variant, clients additionally sort feature importances, adding an $O(T2^{M}d\log d)$ overhead.
  • Server-side computation: The server never touches raw data. It solely aggregates $K$ sketches of size $B$ across $d$ features for up to $2^{M}$ nodes per tree. This results in a server computation complexity of $O(T2^{M}dBK)$.
Table 1: Communication and computation costs required for the entire training phase across centralized, local, and federated methods.
Method Comm.: Client $\to$ Server Comm.: Server $\to$ all Clients Computation: Client Computation: Server
Centralized RF N/A N/A N/A $O(T2^{M}dn)$
FedForest Quantiles $O(T2^{M}dB)$ $O(KT2^{M}dB)$ $O(TMdn_{k})$ $O(T2^{M}dBK)$
FedForest AvgImp Top-$L$ $O(T2^{M}(d+LB))$ $O(KT2^{M}LB)$ $O(TMdn_{k}+T2^{M}d\log d)$ $O(KT2^{M}(d+LB))$
Federated Histograms $O(T2^{M}dB_{\text{hist}})$ $O(KT2^{M})$ $O(TMdn_{k})$ $O(T2^{M}dB_{\text{hist}}K)$
Local Ensembling $O(T2^{M})$ N/A $O(TMdn_{k})$ N/A
Parametric (FedAvg) $O(T_{\text{comm}}P)$ $O(KT_{\text{comm}}P)$ $O(T_{\text{comm}}Pn_{k})$ $O(T_{\text{comm}}KP)$
Table 1 highlights several scalability properties of federated tree training. First, the communication cost of FedForest is independent of the total sample size $n$: once local summaries have been computed, communication depends only on the forest structure $(T,M)$, the feature dimension $d$, and the sketch resolution $B$. Second, Top-$L$ shortlisting reduces the number of feature–threshold pairs exchanged and evaluated. In high-dimensional regimes with $L\ll d$, this substantially lowers the server-to-client broadcast cost, from $O(KT2^{M}dB)$ to $O(KT2^{M}LB)$ for the evaluation stage. Third, the computational load is naturally distributed: clients perform the data-dependent operations on their local samples, while the server only aggregates sketches and sufficient statistics. Unlike FedAvg, which requires an undetermined number of communication rounds ($T_{\text{comm}}$) to converge, FedForest proceeds level by level and therefore requires a deterministic $O(M)$ synchronized rounds for trees of depth at most $M$.
Appendix B Benchmark on Public Datasets
We further assess robustness on 13 public tabular datasets, comprising 6 regression and 7 classification tasks from OpenML and scikit-learn. Since these datasets are not naturally multicentric, we simulate a federated setting by partitioning each dataset across $K=20$ clients using a Dirichlet allocation. The concentration parameter $\alpha$ controls the degree of heterogeneity: small values, such as $\alpha=0.1$, induce highly non-i.i.d. partitions with pronounced covariate shift, while larger values yield increasingly homogeneous client distributions.
We compare FedForest against the tree-based baselines considered in the main text, as well as a federated neural-network baseline with two hidden layers of sizes $128$ and $64$.
Table 2: Performance comparison of centralized, local, and federated learning methods across 14 datasets at varying levels of covariate heterogeneity ($\alpha$). The seven first datasets are regression tasks and performance is evaluated in terms of $R^{2}$. The seven next datasets are classification tasks and evaluation is in terms of balanced accuracy. Top-1 performers are written in bold, top-2 are underlined.
Dataset $\alpha$ Cent. Scikit Cent. 2-Layer NN FedForest Quantiles FedForest AvgImp Fed. Histogram Fed. Ensemble Local Learning
diabetes 10 $0.40\pm 0.06$ $0.36\pm 0.06$ $0.41\pm 0.07$ $0.41\pm 0.05$ $0.41\pm 0.05$ $0.33\pm 0.05$ $0.26\pm 0.09$
diabetes 1 $0.46\pm 0.05$ $0.3\pm 0.3$ $6\pm 0.05 0.44±0$$4\pm 0.06$ $0.44\pm 0.06$ $0.31\pm 0.04$ $0.24\pm 0.08$
diabetes 0.1 $0.43\pm 0.02$ $0.3\pm 0.2$ $3\pm 0.04 0.44±0$$4\pm 0.04$ $0.43\pm 0.05$ $0.21\pm 0.04$ $0.25\pm 0.02$
bodyfat 10 $0.92\pm 0.02$ $0.89\pm 0.06$ $0.93\pm 0.02$ $0.92\pm 0.02$ $0.92\pm 0.02$ $0.25\pm 0.03$ $0.46\pm 0.03$
bodyfat 1 $0.90\pm 0.02$ $0.91\pm 0.03$ $0.91\pm 0.02$ $0.90\pm 0.01$ $0.90\pm 0.02$ $0.26\pm 0.07$ $0.3\pm 0.2$
bodyfat 0.1 $0.90\pm 0.02$ $0.87\pm 0.08$ $0.91\pm 0.02$ $0.89\pm 0.02$ $0.91\pm 0.03$ $0.25\pm 0.06$ $0.6\pm 0.1$
space-ga 10 $0.57\pm 0.04$ $0.71\pm 0.03$ $0.58\pm 0.04$ $0.57\pm 0.04$ $0.59\pm 0.03$ $0.30\pm 0.02$ $0.28\pm 0.02$
space-ga 1 $0.57\pm 0.04$ $0.71\pm 0.03$ $0.58\pm 0.05$ $0.56\pm 0.04$ $0.58\pm 0.04$ $0.28\pm 0.03$ $0.29\pm 0.02$
space-ga 0.1 $0.56\pm 0.03$ $0.71\pm 0.02$ $0.57\pm 0.03$ $0.56\pm 0.02$ $0.57\pm 0.03$ $0.19\pm 0.03$ $0.34\pm 0.02$
puma8NH 10 $0.57\pm 0.01$ $0.67\pm 0.01$ $0.58\pm 0.01$ $0.548\pm 0.007$ $0.569\pm 0.010$ $0.194\pm 0.003$ $0.190\pm 0.004$
puma8NH 1 $0.578\pm 0.007$ $0.678\pm 0.005$ $0.589\pm 0.005$ $0.563\pm 0.006$ $0.578\pm 0.006$ $0.196\pm 0.005$ $0.206\pm 0.008$
puma8NH 0.1 $0.568\pm 0.009$ $0.675\pm 0.008$ $0.581\pm 0.008$ $0.556\pm 0.008$ $0.575\pm 0.005$ $0.155\pm 0.008$ $0.26\pm 0.02$
cpu-act 10 $0.971\pm 0.003$ $0.977\pm 0.002$ $0.962\pm 0.003$ $0.967\pm 0.004$ $0.975\pm 0.003$ $0.29\pm 0.02$ $0.28\pm 0.02$
cpu-act 1 $0.972\pm 0.003$ $0.976\pm 0.003$ $0.956\pm 0.003$ $0.966\pm 0.003$ $0.976\pm 0.002$ $0.29\pm 0.01$ $0.33\pm 0.04$
cpu-act 0.1 $0.973\pm 0.002$ $0.978\pm 0.002$ $0.959\pm 0.001$ $0.970\pm 0.002$ $0.976\pm 0.002$ $0.25\pm 0.02$ $0.47\pm 0.04$
kin8nm 10 $0.496\pm 0.007$ $0.905\pm 0.005$ $0.534\pm 0.002$ $0.501\pm 0.005$ $0.527\pm 0.007$ $0.252\pm 0.003$ $0.245\pm 0.005$
kin8nm 1 $0.504\pm 0.006$ $0.908\pm 0.005$ $0.538\pm 0.006$ $0.506\pm 0.006$ $0.540\pm 0.008$ $0.246\pm 0.006$ $0.253\pm 0.005$
kin8nm 0.1 $0.503\pm 0.007$ $0.904\pm 0.007$ $0.535\pm 0.007$ $0.520\pm 0.008$ $0.536\pm 0.005$ $0.20\pm 0.02$ $0.286\pm 0.009$
breast-cancer 10 $0.951\pm 0.009$ $0.92\pm 0.03$ $0.96\pm 0.02$ $0.96\pm 0.01$ $0.95\pm 0.02$ $0.94\pm 0.01$ $0.90\pm 0.03$
breast-cancer 1 $0.95\pm 0.02$ $0.92\pm 0.05$ $0.95\pm 0.02$ $0.95\pm 0.02$ $0.95\pm 0.01$ $0.91\pm 0.02$ $0.90\pm 0.01$
breast-cancer 0.1 $0.964\pm 0.006$ $0.93\pm 0.03$ $0.960\pm 0.007$ $0.96\pm 0.01$ $0.96\pm 0.01$ $0.90\pm 0.02$ $0.91\pm 0.02$
Spectf 10 $0.81\pm 0.07$ $0.61\pm 0.02$ $0.84\pm 0.08$ $0.79\pm 0.09$ $0.83\pm 0.08$ $0.5000$ $0.61\pm 0.06$
Spectf 1 $0.85\pm 0.03$ $0.63\pm 0.07$ $0.82\pm 0.04$ $0.82\pm 0.03$ $0.83\pm 0.02$ $0.5000$ $0.60\pm 0.03$
Spectf 0.1 $0.81\pm 0.05$ $0.61\pm 0.06$ $0.82\pm 0.07$ $0.78\pm 0.05$ $0.76\pm 0.04$ $0.5000$ $0.63\pm 0.06$
Bioresponse 10 $0.760\pm 0.009$ $71\pm 0.005 0.77±0$$7\pm 0.01$ $0.753\pm 0.009$ $0.76\pm 0.01$ $0.63\pm 0.02$ $0.63\pm 0.01$
Bioresponse 1 $53\pm 0.009 0.77±0$$7\pm 0.01$ $0.76\pm 0.01$ $0.75\pm 0.01$ $0.754\pm 0.008$ $0.638\pm 0.008$ $0.631\pm 0.010$
Bioresponse 0.1 $6\pm 0.01 0.77±0$$7\pm 0.02$ $0.765\pm 0.006$ $0.75\pm 0.01$ $0.763\pm 0.005$ $0.62\pm 0.03$ $0.647\pm 0.008$
madelon 10 $0.68\pm 0.02$ $0.55\pm 0.01$ $0.679\pm 0.007$ $0.64\pm 0.02$ $0.67\pm 0.02$ $0.59\pm 0.01$ $0.56\pm 0.02$
madelon 1 $0.68\pm 0.02$ $0.54\pm 0.01$ $0.69\pm 0.01$ $3\pm 0.01 0.66±0$$6\pm 0.02$ $0.59\pm 0.03$ $0.569\pm 0.008$
madelon 0.1 $0.67\pm 0.02$ $0.560\pm 0.009$ $84\pm 0.010 0.66±0$$6\pm 0.02$ $0.66\pm 0.01$ $0.58\pm 0.05$ $0.61\pm 0.02$
Wine 10 $0.98\pm 0.02$ $0.8\pm 0.1$ $0.990\pm 0.009$ $0.98\pm 0.02$ $0.98\pm 0.02$ $0.3333$ $0.86\pm 0.06$
Wine 1 $0.97\pm 0.02$ $0.81\pm 0.10$ $0.99\pm 0.01$ $0.97\pm 0.02$ $0.97\pm 0.02$ $0.3333$ $0.75\pm 0.06$
Wine 0.1 $0.98\pm 0.02$ $0.8\pm 0.1$ $0.98\pm 0.02$ $80\pm 0.007 0.990$$90\pm 0.009$ $0.3333$ $0.81\pm 0.04$
Bank-Note 10 $0.991\pm 0.007$ $0.97\pm 0.02$ $0.992\pm 0.007$ $0.989\pm 0.006$ $0.992\pm 0.006$ $0.91\pm 0.02$ $0.891\pm 0.006$
Bank-Note 1 $0.988\pm 0.006$ $0.97\pm 0.01$ $0.989\pm 0.004$ $0.989\pm 0.003$ $0.987\pm 0.003$ $9\pm 0.02 0.88±0$$8\pm 0.02$
Bank-Note 0.1 $0.990\pm 0.004$ $0.980\pm 0.007$ $92\pm 0.003 0.990$$90\pm 0.004$ $0.989\pm 0.003$ $0.92\pm 0.02$ $0.92\pm 0.02$
Wisconsin 10 $0.959\pm 0.007$ $0.93\pm 0.04$ $0.968\pm 0.008$ $0.96\pm 0.01$ $0.968\pm 0.007$ $0.937\pm 0.009$ $0.93\pm 0.02$
Wisconsin 1 $0.96\pm 0.01$ $0.95\pm 0.02$ $0.972\pm 0.006$ $0.972\pm 0.006$ $0.964\pm 0.008$ $0.930\pm 0.008$ $0.91\pm 0.02$
Wisconsin 0.1 $0.96\pm 0.01$ $0.96\pm 0.02$ $0.97\pm 0.01$ $0.97\pm 0.01$ $0.97\pm 0.01$ $0.87\pm 0.05$ $0.93\pm 0.03$
As $\alpha$ decreases, client distributions become more fragmented and local methods degrade accordingly. In contrast, FedForest Quantiles and FedForest AvgImp remain stable under covariate shift and closely track the centralized scikit-learn Random Forest baseline. This supports our claim that aggregating split information across clients is substantially more effective than training isolated local trees, even when those local predictors are later ensembled.
Across the benchmark, the proposed FedForest variants consistently rank among the best-performing methods. FedForest Quantiles is typically a top performer across tasks and heterogeneity regimes, while FedForest AvgImp achieves comparable performance in many cases. This suggests that averaging local impurity gains is an effective practical approximation, despite its stronger theoretical guarantees being restricted to homogeneous settings.
Tree-based federated methods also compare favorably with the neural-network baseline, especially in more heterogeneous regimes where flexible nonparametric splits appear better adapted to the data structure. Federated Histograms remain competitive, but FedForest Quantiles often matches or improves upon them, supporting the use of quantile-based candidates rather than fixed-width bins.
Performance in Highly Fragmented Settings
To further evaluate the robustness of our approaches in highly fragmented scenarios, we conducted additional experiments on the Banknote Authentication dataset. We varied the number of clients from $K=1$ (centralized) up to $K=100$, maintaining the same base experimental setup.
As shown in Table 3, FedForest remains highly stable as fragmentation increases. The performance of both FedForest variants stays remarkably close to the Centralized Scikit RF baseline across all values of $K$. In stark contrast, Local Learning and Federated Ensemble approaches degrade sharply as the data becomes divided into smaller, isolated client pools.
These results demonstrate that the primary difficulty in fragmented settings is not the quantile construction itself, but rather the inherent failure of purely local approaches to recover the global split structure when individual clients lack sufficient data.
Table 3: Predictive performance on the Banknote Authentication dataset across varying degrees of client fragmentation ($K=1$ to $K=100$). Values indicate mean $\pm$ standard deviation.
Method K=1 K=2 K=5 K=10 K=20 K=50 K=100
Centralized Scikit RF $0.985\pm 0.007$ $0.991\pm 0.002$ $0.986\pm 0.009$ $0.987\pm 0.005$ $91\pm 0.006 0.990$$90\pm 0.004$ $0.991\pm 0.001$
Fed. Ensemble (Hauschild et al.) $0.988\pm 0.005$ $0.986\pm 0.005$ $0.96\pm 0.01$ $0.93\pm 0.01$ $0.91\pm 0.01$ $0.87\pm 0.02$ $0.5000$
Fed. Histogram (Kalloori et al.) $0.996\pm 0.001$ $0.998\pm 0.004$ $0.996\pm 0.003$ $0.994\pm 0.004$ $0.994\pm 0.005$ $0.992\pm 0.007$ $0.997\pm 0.001$
FedForest AvgImp (ours) $0.994\pm 0.002$ $0.996\pm 0.004$ $0.993\pm 0.006$ $0.991\pm 0.004$ $0.996\pm 0.001$ $0.988\pm 0.003$ $0.989\pm 0.003$
FedForest Quantiles (ours) $0.988\pm 0.008$ $0.993\pm 0.001$ $86\pm 0.007 0.990$$90\pm 0.005$ $0.9953\pm 0.0006$ $0.991\pm 0.007$ $0.993\pm 0.002$
Local Learning $0.988\pm 0.005$ $0.980\pm 0.005$ $0.949\pm 0.005$ $0.913\pm 0.007$ $0.89\pm 0.01$ $0.79\pm 0.03$ $0.55\pm 0.05$
Appendix C Proofs
C.1 Candidate thresholds with quantile sketching
Proof of Theorem 3.1.